Thompson Aeronautics repairs aircraft engines. The company’s Purchasing Department supports its two departments, Defense and Commercial. The Defense division has contracts with the Department of Defense and the Commercial division works primarily with domestic airlines and air freight companies. The cost of the Purchasing Department is $6.70 million annually. Information on the activity of the Purchasing Department for the last year follows: Number of Purchase Orders Dollar Amount of Purchases Defense 8,800 $ 132,000,000 Commercial 35,200 198,000,000

Required: a. What is the cost charged to each division if Thompson allocates Purchasing Department costs based on the number of purchase orders? (Do not round intermediate calculations. Enter your answers in thousands of dollars.)

b. What is the cost charged to each division if Thompson allocates Purchasing Department costs based on the dollar amount of the purchases? (Do not round intermediate calculations. Enter your answers in thousands of dollars.)
Answer #1
a
Cost Allocation
Defense 1340000 or 1340 =6700000/(8800+35200)*8800
Commercial 5360000 or 5360 =6700000/(8800+35200)*35200
b
Cost Allocation
Defense 2680000 or 2680 =6700000/(132+198)*132
Commercial 4020000 or 4020 =6700000/(132+198)*198